\begin{thebibliography}{10}

\bibitem{1_SPIE22_davies1983quality}
Davies, E.~R., ``Quality assurance in diagnostic radiology, world health
  organisation, eastbourne (1982) quality assurance in nuclear medicine, world
  health organisation, switzerland (1982), 70 pp,'' (1982).

\bibitem{12_wismuller2006exploratory}
Wism{\"u}ller, A., ``Exploratory morphogenesis ({XOM}): a novel computational
  framework for self-organization,'' {\em Ph. D. thesis, Technical University
  of Munich, Department of Electrical and Computer Engineering}  (2006).

\bibitem{13_wismuller1998neural}
Wism{\"u}ller, A., Dersch, D.~R., Lipinski, B., Hahn, K., and Auer, D., ``A
  neural network approach to functional {MRI} pattern analysis—clustering of
  time-series by hierarchical vector quantization,'' in [{\em International
  Conference on Artificial Neural Networks}{\nolinebreak\hspace{0.1em}]},
  857--862, Springer (1998).

\bibitem{14_wismuller2002deformable}
Wism{\"u}ller, A., Vietze, F., Dersch, D.~R., Behrends, J., Hahn, K., and
  Ritter, H., ``The deformable feature map-a novel neurocomputing algorithm for
  adaptive plasticity in pattern analysis,'' {\em Neurocomputing}~{\bf
  48}(1-4),  107--139 (2002).

\bibitem{15_behrends2003segmentation}
Behrends, J., Hoole, P., Leinsinger, G.~L., Tillmann, H.~G., Hahn, K., Reiser,
  M., and Wism{\"u}ller, A., ``A segmentation and analysis method for {MRI}
  data of the human vocal tract,'' in [{\em Bildverarbeitung f{\"u}r die
  Medizin 2003}{\nolinebreak\hspace{0.1em}]},   186--190, Springer (2003).

\bibitem{16_wismuller1997neural}
Wism{\"u}ller, A., ``Neural network computation in biomedical research: chances
  for conceptual cross-fertilization,'' {\em Theory in Biosciences}  (1997).

\bibitem{17_bunte2010exploratory}
Bunte, K., Hammer, B., Villmann, T., Biehl, M., and Wism{\"u}ller, A.,
  ``Exploratory observation machine ({XOM}) with {K}ullback-{L}eibler
  divergence for dimensionality reduction and visualization.,'' in [{\em
  ESANN}{\nolinebreak\hspace{0.1em}]},   {\bf 10},  87--92 (2010).

\bibitem{18_wismuller1998deformable}
Wism{\"u}ller, A., Vietze, F., Dersch, D.~R., Hahn, K., and Ritter, H., ``The
  deformable feature map—adaptive plasticity for function approximation,'' in
  [{\em International Conference on Artificial Neural
  Networks}{\nolinebreak\hspace{0.1em}]},   123--128, Springer (1998).

\bibitem{19_wismuller2009exploration}
Wism{\"u}ller, A., ``The exploration machine--a novel method for data
  visualization,'' in [{\em International Workshop on Self-Organizing
  Maps}{\nolinebreak\hspace{0.1em}]},   344--352, Springer (2009).

\bibitem{20_wismuller2009method}
Wism{\"u}ller, A., ``Method, data processing device and computer program
  product for processing data,'' (July~28 2009).
\newblock US Patent 7,567,889.

\bibitem{22_huber2010classification}
Huber, M.~B., Nagarajan, M., Leinsinger, G., Ray, L.~A., and Wism{\"u}ller, A.,
  ``Classification of interstitial lung disease patterns with topological
  texture features,'' in [{\em Medical Imaging 2010: Computer-Aided
  Diagnosis}{\nolinebreak\hspace{0.1em}]},   {\bf 7624},  762410, International
  Society for Optics and Photonics (2010).

\bibitem{23_wismuller2009exploration}
Wism{\"u}ller, A., ``The exploration machine: a novel method for analyzing
  high-dimensional data in computer-aided diagnosis,'' in [{\em Medical Imaging
  2009: Computer-Aided Diagnosis}{\nolinebreak\hspace{0.1em}]},   {\bf 7260},
  72600G, International Society for Optics and Photonics (2009).

\bibitem{24_bunte2011neighbor}
Bunte, K., Hammer, B., Villmann, T., Biehl, M., and Wism{\"u}ller, A.,
  ``Neighbor embedding {XOM} for dimension reduction and visualization,'' {\em
  Neurocomputing}~{\bf 74}(9),  1340--1350 (2011).

\bibitem{25_meyer2004model}
Meyer-B{\"a}se, A., Lange, O., Wism{\"u}ller, A., and Ritter, H., ``Model-free
  functional {MRI} analysis using topographic independent component analysis,''
  {\em International journal of neural systems}~{\bf 14}(04),  217--228 (2004).

\bibitem{wismuller2001exploration}
Wism{\"u}ller, A., ``Exploration-organized morphogenesis ({XOM}): A general
  framework for learning by self-organization,'' {\em
  Forschungsberichte-Institut f{\"u}r Phonetik und Sprachliche Kommunikation
  der Universit{\"a}t M{\"u}nchen} (37),  205--239 (2001).

\bibitem{saalbach2005hyperbolic}
Saalbach, A., Twellmann, T., Nattkemper, T.~W., Wism{\"u}ller, A., Ontrup, J.,
  and Ritter, H.~J., ``A hyperbolic topographic mapping for proximity data.,''
  in [{\em Artificial Intelligence and
  Applications}{\nolinebreak\hspace{0.1em}]},   {\bf 2005},  106--111 (2005).

\bibitem{wismuller2009computational}
Wism{\"u}ller, A., ``A computational framework for exploratory data
  analysis.,'' in [{\em ESANN}{\nolinebreak\hspace{0.1em}]},  Citeseer (2009).

\bibitem{26_wismuller2009computational}
Wism{\"u}ller, A., ``A computational framework for nonlinear dimensionality
  reduction and clustering,'' in [{\em International Workshop on
  Self-Organizing Maps}{\nolinebreak\hspace{0.1em}]},   334--343, Springer
  (2009).

\bibitem{leinsinger2003volumetric}
Leinsinger, G., Teipel, S., Wism{\"u}ller, A., Born, C., Meindl, T., Flatz, W.,
  Sch{\"o}nberg, S., Pruessner, J., Hampel, H., and Reiser, M., ``Volumetric
  {MRI} for evaluation of regional pattern and progressin of neocortical
  degeneration in alzheimer's disease,'' {\em Der Radiologe}~{\bf 43}(7),
  537--542 (2003).

\bibitem{27_meyer2003topographic}
Meyer-Base, A., Auer, D., and Wism{\"u}ller, A., ``Topographic independent
  component analysis for f{MRI} signal detection,'' in [{\em Proceedings of the
  International Joint Conference on Neural Networks,
  2003.}{\nolinebreak\hspace{0.1em}]},   {\bf 1},  601--605, IEEE (2003).

\bibitem{wismuller2009exploration}
Wism{\"u}ller, A., ``The exploration machine-a novel method for
  structure-preserving dimensionality reduction.,'' in [{\em
  ESANN}{\nolinebreak\hspace{0.1em}]},  Citeseer (2009).

\bibitem{28_meyer2009small}
Meyer-Baese, A., Schlossbauer, T., Lange, O., and Wism{\"u}ller, A., ``Small
  lesions evaluation based on unsupervised cluster analysis of signal-intensity
  time courses in dynamic breast {MRI},'' {\em International journal of
  biomedical imaging}~{\bf 2009} (2009).

\bibitem{29_wismueller2010model}
Wism{\"u}ller, A., Lange, O., Auer, D., and Leinsinger, G., ``Model-free
  functional {MRI} analysis for detecting low-frequency functional connectivity
  in the human brain,'' in [{\em Medical Imaging 2010: Computer-Aided
  Diagnosis}{\nolinebreak\hspace{0.1em}]},   {\bf 7624},  76241M, International
  Society for Optics and Photonics (2010).

\bibitem{meyer2007unsupervised}
Meyer-B{\"a}se, A., Saalbach, A., Lange, O., and Wism{\"u}ller, A.,
  ``Unsupervised clustering of f{MRI} and {MRI} time series,'' {\em Biomedical
  Signal Processing and Control}~{\bf 2}(4),  295--310 (2007).

\bibitem{wismuller2000neural}
Wism{\"u}ller, A., Vietze, F., Dersch, D.~R., Hahn, K., and Ritter, H.~J., ``A
  neural network approach to adaptive pattern analysis-the deformable feature
  map.,'' in [{\em ESANN}{\nolinebreak\hspace{0.1em}]},   189--194 (2000).

\bibitem{meyer2007analysis}
Meyer-Baese, A., Lange, O., Wism{\"u}ller, A., and Hurdal, M.~K., ``Analysis of
  dynamic susceptibility contrast {MRI} time series based on unsupervised
  clustering methods,'' {\em IEEE Transactions on Information Technology in
  Biomedicine}~{\bf 11}(5),  563--573 (2007).

\bibitem{32_wismueller2008human}
Wism{\"u}ller, A., Behrends, J., Hoole, P., Leinsinger, G.~L., Reiser, M.~F.,
  and Westesson, P.-L., ``Human vocal tract analysis by in vivo 3d {MRI} during
  phonation: a complete system for imaging, quantitative modeling, and speech
  synthesis,'' in [{\em International Conference on Medical Image Computing and
  Computer-Assisted Intervention}{\nolinebreak\hspace{0.1em}]},   306--312,
  Springer (2008).

\bibitem{wismuller2000hierarchical}
Wism{\"u}ller, A., Dersch, D.~R., Lipinski, B., Hahn, K., and Auer, D.,
  ``Hierarchical clustering of functional {MRI} time-series by deterministic
  annealing,'' in [{\em International Symposium on Medical Data
  Analysis}{\nolinebreak\hspace{0.1em}]},   49--54, Springer (2000).

\bibitem{wismuller2015method}
Wism{\"u}ller, A., ``Method and device for representing multichannel image
  data,'' (Nov.~17 2015).
\newblock US Patent 9,189,846.

\bibitem{33_huber2012texture}
Huber, M.~B., Bunte, K., Nagarajan, M.~B., Biehl, M., Ray, L.~A., and
  Wism{\"u}ller, A., ``Texture feature ranking with relevance learning to
  classify interstitial lung disease patterns,'' {\em Artificial intelligence
  in medicine}~{\bf 56}(2),  91--97 (2012).

\bibitem{37_otto2003model}
Otto, T.~D., Meyer-Baese, A., Hurdal, M., Sumners, D., Auer, D., and
  Wism{\"u}ller, A., ``Model-free functional {MRI} analysis using cluster-based
  methods,'' in [{\em Intelligent Computing: Theory and
  Applications}{\nolinebreak\hspace{0.1em}]},   {\bf 5103},  17--24,
  International Society for Optics and Photonics (2003).

\bibitem{38_varini2004breast}
Varini, C., Nattkemper, T.~W., Degenhard, A., and Wism{\"u}ller, A., ``Breast
  {MRI} data analysis by lle,'' in [{\em 2004 IEEE International Joint
  Conference on Neural Networks (IEEE Cat. No.
  04CH37541)}{\nolinebreak\hspace{0.1em}]},   {\bf 3},  2449--2454, IEEE
  (2004).

\bibitem{48_meyer2004computer}
Meyer-Baese, A., Wism{\"u}ller, A., Lange, O., and Leinsinger, G.,
  ``Computer-aided diagnosis in breast {MRI} based on unsupervised clustering
  techniques,'' in [{\em Intelligent Computing: Theory and Applications
  II}{\nolinebreak\hspace{0.1em}]},   {\bf 5421},  29--37, International
  Society for Optics and Photonics (2004).

\bibitem{40_meyer2004stability}
Meyer-Base, A., Pilyugin, S.~S., and Wism{\"u}ller, A., ``Stability analysis of
  a self-organizing neural network with feedforward and feedback dynamics,'' in
  [{\em 2004 IEEE International Joint Conference on Neural Networks (IEEE Cat.
  No. 04CH37541)}{\nolinebreak\hspace{0.1em}]},   {\bf 2},  1505--1509, IEEE
  (2004).

\bibitem{wismuller1998hierarchical}
Wism{\"u}ller, A., Dersch, D., Lipinski, B., Hahn, K., and Auer, D.,
  ``Hierarchical clustering of {fMRI} time-series by deterministic annealing,''
  {\em Neuroimage}~{\bf 7}(4),  S593 (1998).

\bibitem{41_meyer2008computer}
Meyer-Baese, A., Lange, O., Schlossbauer, T., and Wism{\"u}ller, A.,
  ``Computer-aided diagnosis and visualization based on clustering and
  independent component analysis for breast {MRI},'' in [{\em 2008 15th IEEE
  International Conference on Image Processing}{\nolinebreak\hspace{0.1em}]},
  3000--3003, IEEE (2008).

\bibitem{wismuller2013introducing}
Wism{\"u}ller, A., De, T., Lochm{\"u}ller, E., Eckstein, F., and Nagarajan,
  M.~B., ``Introducing anisotropic minkowski functionals and quantitative
  anisotropy measures for local structure analysis in biomedical imaging,'' in
  [{\em Medical Imaging 2013: Biomedical Applications in Molecular, Structural,
  and Functional Imaging}{\nolinebreak\hspace{0.1em}]},   {\bf 8672},  86720I,
  International Society for Optics and Photonics (2013).

\bibitem{45_bhole20143d}
Bhole, C., Pal, C., Rim, D., and Wism{\"u}ller, A., ``3d segmentation of
  abdominal ct imagery with graphical models, conditional random fields and
  learning,'' {\em Machine vision and applications}~{\bf 25}(2),  301--325
  (2014).

\bibitem{46_nagarajan2013computer}
Nagarajan, M.~B., Coan, P., Huber, M.~B., Diemoz, P.~C., Glaser, C., and
  Wism{\"u}ller, A., ``Computer-aided diagnosis in phase contrast imaging x-ray
  computed tomography for quantitative characterization of ex vivo human
  patellar cartilage,'' {\em IEEE Transactions on Biomedical Engineering}~{\bf
  60}(10),  2896--2903 (2013).

\bibitem{wismuller2001automatic}
Wism{\"u}ller, A., Behrends, J., Dersch, D., Leinsinger, G., Vietze, F., and
  Hahn, K., ``Automatic segmentation of cerebral contours in multispectral
  {MRI} data sets of the human brain by self-organizing neural networks,'' in
  [{\em Radiology}{\nolinebreak\hspace{0.1em}]},   {\bf 221},  461--461 (2001).

\bibitem{wismueller1999adaptive}
Wismueller, A., Vietze, F., Dersch, D., Leinsinger, G., Ritter, H., and Hahn,
  K., ``Adaptive self-organized template matching of the gray-level feature
  space for automatic segmentation of multispectral {MRI} data of the human
  brain,'' in [{\em Radiology}{\nolinebreak\hspace{0.1em}]},   {\bf 213},
  364--364 (1999).

\bibitem{meyer2005computer}
Meyer-Baese, A., Lange, O., Wismuller, A., and Leinsinger, G., ``Computer-aided
  diagnosis in breast {MRI} based on {ICA} and unsupervised clustering
  techniques,'' in [{\em Independent Component Analyses, Wavelets, Unsupervised
  Smart Sensors, and Neural Networks III}{\nolinebreak\hspace{0.1em}]},   {\bf
  5818},  38--49, International Society for Optics and Photonics (2005).

\bibitem{49_nagarajan2014computer}
Nagarajan, M.~B., Coan, P., Huber, M.~B., Diemoz, P.~C., Glaser, C., and
  Wism{\"u}ller, A., ``Computer-aided diagnosis for phase-contrast x-ray
  computed tomography: quantitative characterization of human patellar
  cartilage with high-dimensional geometric features,'' {\em Journal of digital
  imaging}~{\bf 27}(1),  98--107 (2014).

\bibitem{pester2013exploring}
Pester, B., Leistritz, L., Witte, H., and Wismueller, A., ``Exploring effective
  connectivity by a {G}ranger causality approach with embedded dimension
  reduction,'' {\em Biomedical Engineering/Biomedizinische Technik}~{\bf
  58}(SI-1-Track-G),  000010151520134172 (2013).

\bibitem{50_nagarajan2014classification}
Nagarajan, M.~B., Huber, M.~B., Schlossbauer, T., Leinsinger, G., Krol, A., and
  Wism{\"u}ller, A., ``Classification of small lesions on dynamic breast {MRI}:
  Integrating dimension reduction and out-of-sample extension into cadx
  methodology,'' {\em Artificial intelligence in medicine}~{\bf 60}(1),  65--77
  (2014).

\bibitem{yang2014improving}
Yang, C.-C., Nagarajan, M.~B., Huber, M.~B., Carballido-Gamio, J., Bauer,
  J.~S., Baum, T.~H., Eckstein, F., Lochm{\"u}ller, E.-M., Majumdar, S., Link,
  T.~M., et~al., ``Improving bone strength prediction in human proximal femur
  specimens through geometrical characterization of trabecular bone
  microarchitecture and support vector regression,'' {\em Journal of electronic
  imaging}~{\bf 23}(1),  013013 (2014).

\bibitem{wismuller2014pair}
Wism{\"u}ller, A., Nagarajan, M.~B., Witte, H., Pester, B., and Leistritz, L.,
  ``Pair-wise clustering of large scale {G}ranger causality index matrices for
  revealing communities,'' in [{\em Medical Imaging 2014: Biomedical
  Applications in Molecular, Structural, and Functional
  Imaging}{\nolinebreak\hspace{0.1em}]},   {\bf 9038},  90381R, International
  Society for Optics and Photonics (2014).

\bibitem{wang2014investigating}
Wang, X., Nagarajan, M.~B., Conover, D., Ning, R., O'Connell, A., and
  Wism{\"u}ller, A., ``Investigating the use of texture features for analysis
  of breast lesions on contrast-enhanced cone beam {CT},'' in [{\em Medical
  Imaging 2014: Biomedical Applications in Molecular, Structural, and
  Functional Imaging}{\nolinebreak\hspace{0.1em}]},   {\bf 9038},  903822,
  International Society for Optics and Photonics (2014).

\bibitem{51_wismuller2014framework}
Wism{\"u}ller, A., Wang, X., DSouza, A.~M., and Nagarajan, M.~B., ``A framework
  for exploring non-linear functional connectivity and causality in the human
  brain: mutual connectivity analysis (mca) of resting-state functional {MRI}
  with convergent cross-mapping and non-metric clustering,'' {\em arXiv
  preprint arXiv:1407.3809}  (2014).

\bibitem{meyer2004local}
Meyer-B{\"a}se, A., Pilyugin, S., Wism{\"u}ller, A., and Foo, S., ``Local
  exponential stability of competitive neural networks with different time
  scales,'' {\em Engineering Applications of Artificial Intelligence}~{\bf
  17}(3),  227--232 (2004).

\bibitem{schmidt2014impact}
Schmidt, C., Pester, B., Nagarajan, M., Witte, H., Leistritz, L., and
  Wism{\"u}ller, A., ``Impact of multivariate {G}ranger causality analyses with
  embedded dimension reduction on network modules,'' in [{\em 2014 36th Annual
  International Conference of the IEEE Engineering in Medicine and Biology
  Society}{\nolinebreak\hspace{0.1em}]},   2797--2800, IEEE (2014).

\bibitem{nagarajan2015integrating}
Nagarajan, M.~B., Coan, P., Huber, M.~B., Diemoz, P.~C., and Wism{\"u}ller, A.,
  ``Integrating dimension reduction and out-of-sample extension in automated
  classification of ex vivo human patellar cartilage on phase contrast x-ray
  computed tomography,'' {\em Plos one}~{\bf 10}(2),  e0117157 (2015).

\bibitem{wismuller2015nonlinear}
Wism{\"u}ller, A., Abidin, A.~Z., D'Souza, A.~M., Wang, X., Hobbs, S.~K.,
  Leistritz, L., and Nagarajan, M.~B., ``Nonlinear functional connectivity
  network recovery in the human brain with mutual connectivity analysis
  {(MCA)}: convergent cross-mapping and non-metric clustering,'' in [{\em
  Medical Imaging 2015: Biomedical Applications in Molecular, Structural, and
  Functional Imaging}{\nolinebreak\hspace{0.1em}]},   {\bf 9417},  94170M,
  International Society for Optics and Photonics (2015).

\bibitem{nagarajan2015characterizing}
Nagarajan, M.~B., Checefsky, W.~A., Abidin, A.~Z., Tsai, H., Wang, X., Hobbs,
  S.~K., Bauer, J.~S., Baum, T., and Wism{\"u}ller, A., ``Characterizing
  trabecular bone structure for assessing vertebral fracture risk on volumetric
  quantitative computed tomography,'' in [{\em Medical Imaging 2015: Biomedical
  Applications in Molecular, Structural, and Functional
  Imaging}{\nolinebreak\hspace{0.1em}]},   {\bf 9417},  94171E, International
  Society for Optics and Photonics (2015).

\bibitem{abidin2015volumetric}
Abidin, A.~Z., Nagarajan, M.~B., Checefsky, W.~A., Coan, P., Diemoz, P.~C.,
  Hobbs, S.~K., Huber, M.~B., and Wism{\"u}ller, A., ``Volumetric
  characterization of human patellar cartilage matrix on phase contrast x-ray
  computed tomography,'' in [{\em Medical Imaging 2015: Biomedical Applications
  in Molecular, Structural, and Functional
  Imaging}{\nolinebreak\hspace{0.1em}]},   {\bf 9417},  94171F, International
  Society for Optics and Photonics (2015).

\bibitem{wismuller2016mutual}
Wism{\"u}ller, A., Abidin, A.~Z., DSouza, A.~M., and Nagarajan, M.~B., ``Mutual
  connectivity analysis {(MCA)} for nonlinear functional connectivity network
  recovery in the human brain using convergent cross-mapping and non-metric
  clustering,'' in [{\em Advances in Self-Organizing Maps and Learning Vector
  Quantization}{\nolinebreak\hspace{0.1em}]},   217--226, Springer (2016).

\bibitem{abidin2016investigating}
Abidin, A.~Z., D'Souza, A.~M., Nagarajan, M.~B., and Wism{\"u}ller, A.,
  ``Investigating changes in brain network properties in hiv-associated
  neurocognitive disease ({HAND}) using mutual connectivity analysis ({MCA}),''
  in [{\em Medical Imaging 2016: Biomedical Applications in Molecular,
  Structural, and Functional Imaging}{\nolinebreak\hspace{0.1em}]},   {\bf
  9788},  97881W, International Society for Optics and Photonics (2016).

\bibitem{52_schmidt2016multivariate}
Schmidt, C., Pester, B., Schmid-Hertel, N., Witte, H., Wism{\"u}ller, A., and
  Leistritz, L., ``A multivariate {G}ranger causality concept towards full
  brain functional connectivity,'' {\em PloS one}~{\bf 11}(4) (2016).

\bibitem{abidin2017classification}
Abidin, A.~Z., Jameson, J., Molthen, R., and Wism{\"u}ller, A.,
  ``Classification of micro-{CT} images using {3D} characterization of bone
  canal patterns in human osteogenesis imperfecta,'' in [{\em Medical Imaging
  2017: Computer-Aided Diagnosis}{\nolinebreak\hspace{0.1em}]},   {\bf 10134},
  1013413, International Society for Optics and Photonics (2017).

\bibitem{abidin2017using}
Abidin, A.~Z., Chockanathan, U., DSouza, A.~M., Inglese, M., and Wism{\"u}ller,
  A., ``Using large-scale {G}ranger causality to study changes in brain network
  properties in the clinically isolated syndrome {(CIS)} stage of multiple
  sclerosis,'' in [{\em Medical Imaging 2017: Biomedical Applications in
  Molecular, Structural, and Functional Imaging}{\nolinebreak\hspace{0.1em}]},
   {\bf 10137},  101371B, International Society for Optics and Photonics
  (2017).

\bibitem{61_dsouza2017exploring}
DSouza, A.~M., Abidin, A.~Z., Leistritz, L., and Wism{\"u}ller, A., ``Exploring
  connectivity with large-scale {G}ranger causality on resting-state functional
  {MRI},'' {\em Journal of neuroscience methods}~{\bf 287},  68--79 (2017).

\bibitem{55_abidin2018deep}
Abidin, A.~Z., Deng, B., DSouza, A.~M., Nagarajan, M.~B., Coan, P., and
  Wism{\"u}ller, A., ``Deep transfer learning for characterizing chondrocyte
  patterns in phase contrast x-ray computed tomography images of the human
  patellar cartilage,'' {\em Computers in biology and medicine}~{\bf 95},
  24--33 (2018).

\bibitem{chockanathan2018resilient}
Chockanathan, U., Abidin, A.~Z., DSouza, A.~M., Schifitto, G., and
  Wism{\"u}ller, A., ``Resilient modular small-world directed brain networks in
  healthy subjects with large-scale {G}ranger causality analysis of
  resting-state functional {MRI},'' in [{\em Medical Imaging 2018: Biomedical
  Applications in Molecular, Structural, and Functional
  Imaging}{\nolinebreak\hspace{0.1em}]},   {\bf 10578},  105780B, International
  Society for Optics and Photonics (2018).

\bibitem{dsouza2018mutual}
DSouza, A.~M., Abidin, A.~Z., Chockanathan, U., Schifitto, G., and
  Wism{\"u}ller, A., ``Mutual connectivity analysis of resting-state functional
  {MRI} data with local models,'' {\em NeuroImage}~{\bf 178},  210--223 (2018).

\bibitem{abidin2019investigating}
Abidin, A.~Z., Dar, I., D'Souza, A.~M., Lin, E.~P., and Wism{\"u}ller, A.,
  ``Investigating a quantitative radiomics approach for brain tumor
  classification,'' in [{\em Medical Imaging 2019: Biomedical Applications in
  Molecular, Structural, and Functional Imaging}{\nolinebreak\hspace{0.1em}]},
   {\bf 10953},  109530B, International Society for Optics and Photonics
  (2019).

\bibitem{abidin2020detecting}
Abidin, A.~Z., DSouza, A.~M., Schifitto, G., and Wism{\"u}ller, A., ``Detecting
  cognitive impairment in {HIV}-infected individuals using mutual connectivity
  analysis of resting state functional {MRI},'' {\em Journal of
  neurovirology}~{\bf 26}(2),  188--200 (2020).

\bibitem{wismuller2020prospective}
Wism{\"u}ller, A. and Stockmaster, L., ``A prospective randomized clinical
  trial for measuring radiology study reporting time on artificial
  intelligence-based detection of intracranial hemorrhage in emergent care head
  ct,'' in [{\em Medical Imaging 2020: Biomedical Applications in Molecular,
  Structural, and Functional Imaging}{\nolinebreak\hspace{0.1em}]},   {\bf
  11317},  113170M, International Society for Optics and Photonics (2020).

\bibitem{dsouza2020large}
Wism{\"u}ller, A., DSouza, A.~M., Abidin, A.~Z., and Vosoughi, M.~A.,
  ``Large-scale nonlinear {G}ranger causality: A data-driven, multivariate
  approach to recovering directed networks from short time-series data,'' {\em
  arXiv preprint arXiv:2009.04681}  (2020).

\bibitem{vosoughi2020large}
Vosoughi, M.~A. and Wism{\"u}ller, A., ``Large-scale kernelized {G}ranger
  causality to infer topology of directed graphs with applications to brain
  networks,'' {\em arXiv preprint arXiv:2011.08261}  (2020).

\bibitem{vosoughi2021marijuana}
Vosoughi, M.~A. and Wism{\"u}ller, A., ``Large-scale extended granger causality
  for classification of marijuana users from functional mri,'' in [{\em Medical
  Imaging 2021: Biomedical Applications in Molecular, Structural, and
  Functional Imaging}{\nolinebreak\hspace{0.1em}]},   {\bf 11600},  116000D,
  International Society for Optics and Photonics (2021).

\bibitem{vosoughi2021eusipco}
Vosoughi, M.~A. and Wism{\"u}ller, A., ``Leveraging pre-images to discover
  nonlinear relationships in multivariate environments,'' in [{\em 2021
  {European Signal Processing Conference
  ({EUSIPCO})}}{\nolinebreak\hspace{0.1em}]},  IEEE (August 2021).

\bibitem{dsouza2021large}
Wism{\"u}ller, A., Dsouza, A.~M., Vosoughi, M.~A., and Abidin, A.,
  ``Large-scale nonlinear {G}ranger causality for inferring directed dependence
  from short multivariate time-series data,'' {\em Scientific reports}~{\bf
  11}(1),  1--11 (2021).

\bibitem{vosoughi2021_LSAGC}
Wism{\"u}ller, A. and Vosoughi, M.~A., ``Large-scale augmented {G}ranger
  causality ({lsAGC}) for connectivity analysis in complex systems: From
  computer simulations to functional {MRI} ({fMRI}),'' in [{\em Medical Imaging
  2021: Biomedical Applications in Molecular, Structural, and Functional
  Imaging}{\nolinebreak\hspace{0.1em}]},   {\bf 11600},  1160009, International
  Society for Optics and Photonics (2021).

\bibitem{vosoughi2021_lsNonlinear}
Wism{\"u}ller, A., DSouza, A.~M., Vosoughi, M.~A., and Z.~Abidin, A., ``Network
  connectivity analysis in complex systems using large-scale non-linear
  {G}ranger causality {(lsNGC)},'' {\em International Society for Optics and
  Photonics}  (2021).

\bibitem{vosoughi2021schizophrenia}
Wism{\"u}ller, A. and Vosoughi, M.~A., ``Classification of schizophrenia from
  functional {MRI} using large-scale extended {G}ranger causality,'' {\em
  International Society for Optics and Photonics}  (2021).

\bibitem{2_SPIE22}
``{RADPEER}.'' \url{https://www.acr.org/Clinical-Resources/RADPEER}.
\newblock Accessed: 2022-03-07.

\bibitem{3_SPIE22_freudenburg1988perceived}
Freudenburg, W.~R., ``Perceived risk, real risk: Social science and the art of
  probabilistic risk assessment,'' {\em Science}~{\bf 242}(4875),  44--49
  (1988).

\bibitem{4_SPIE22_krause2012vigilance}
Krause, M., ``Vigilance fatigue in policing: A critical threat to public safety
  and officer well-being,'' {\em FBI L. Enforcement Bull.}~{\bf 81},  3 (2012).

\bibitem{5_SPIOE22_powers20182018}
Powers, W.~J., Rabinstein, A.~A., Ackerson, T., Adeoye, O.~M., Bambakidis,
  N.~C., Becker, K., Biller, J., Brown, M., Demaerschalk, B.~M., Hoh, B.,
  et~al., ``2018 guidelines for the early management of patients with acute
  ischemic stroke: a guideline for healthcare professionals from the {American
  Heart Association/American Stroke Association},'' {\em stroke}~{\bf 49}(3),
  e46--e99 (2018).

\bibitem{6_SPIE22_lees2010time}
Lees, K.~R., Bluhmki, E., Von~Kummer, R., Brott, T.~G., Toni, D., Grotta,
  J.~C., Albers, G.~W., Kaste, M., Marler, J.~R., Hamilton, S.~A., et~al.,
  ``Time to treatment with intravenous alteplase and outcome in stroke: an
  updated pooled analysis of ecass, atlantis, ninds, and epithet trials,'' {\em
  The Lancet}~{\bf 375}(9727),  1695--1703 (2010).

\bibitem{7_SPIE22_walter2012diagnosis}
Walter, S., Kostopoulos, P., Haass, A., Keller, I., Lesmeister, M.,
  Schlechtriemen, T., Roth, C., Papanagiotou, P., Grunwald, I., Schumacher, H.,
  et~al., ``Diagnosis and treatment of patients with stroke in a mobile stroke
  unit versus in hospital: a randomised controlled trial,'' {\em The Lancet
  Neurology}~{\bf 11}(5),  397--404 (2012).

\bibitem{8_SPIE22_katzan2004utilization}
Katzan, I.~L., Hammer, M.~D., Hixson, E.~D., Furlan, A.~J., Abou-Chebl, A., and
  Nadzam, D.~M., ``Utilization of intravenous tissue plasminogen activator for
  acute ischemic stroke,'' {\em Archives of neurology}~{\bf 61}(3),  346--350
  (2004).

\bibitem{9_SPIE22_mazighi2013impact}
Mazighi, M., Chaudhry, S.~A., Ribo, M., Khatri, P., Skoloudik, D., Mokin, M.,
  Labreuche, J., Meseguer, E., Yeatts, S.~D., Siddiqui, A.~H., et~al., ``Impact
  of onset-to-reperfusion time on stroke mortality: a collaborative pooled
  analysis,'' {\em Circulation}~{\bf 127}(19),  1980--1985 (2013).

\bibitem{10_SPIE22_wismuller2020prospective}
Wism{\"u}ller, A. and Stockmaster, L., ``A prospective randomized clinical
  trial for measuring radiology study reporting time on {Artificial
  Intelligence}-based detection of intracranial hemorrhage in emergent care
  head ct,'' in [{\em Medical Imaging 2020: Biomedical Applications in
  Molecular, Structural, and Functional Imaging}{\nolinebreak\hspace{0.1em}]},
   {\bf 11317},  113170M, International Society for Optics and Photonics
  (2020).

\end{thebibliography}

\documentclass[]{spie}  %>>> use for US letter paper

\renewcommand{\baselinestretch}{1.65} % Change to 1.65 for double spacing
\usepackage{subcaption}
\usepackage{amsmath,amsfonts,amssymb}
\usepackage{graphicx}
\usepackage[colorlinks=true, allcolors=blue]{hyperref}
\usepackage[leftcaption]{sidecap}

\title{Re-defining Radiology Quality Assurance (QA) - Artificial Intelligence 
(AI)-Based QA by Restricted Investigation of Unequal Scores 
(AQUARIUS) }

\author[a,b,c,d]{Axel Wismüller}
\author[a]{Larry Stockmaster}
\author[b]{M. Ali Vosoughi}

\affil[a]{Department of Imaging Sciences, University of Rochester, NY, USA}
\affil[b]{Department of Electrical and Computer Engineering, University of Rochester, NY, USA}
\affil[c]{Department of Biomedical Engineering, University of Rochester, NY, USA}
\affil[d]{Faculty of Medicine and Institute of Clinical Radiology, Ludwig Maximilian University,
Munich, Germany}

\authorinfo{Further author information: (Send correspondence to Axel Wismüller)\\Axel Wismüller: E-mail: axel.wismueller@gmail.com}

\begin{document} 
\maketitle

\begin{abstract}
There is an urgent need for streamlining radiology Quality Assurance (QA) programs to make them better and faster. Here, we present a novel approach, \textbf{A}rtificial Intelligence (AI)-Based \textbf{Qu}ality \textbf{A}ssurance by \textbf{R}estricted \textbf{I}nvestigation of \textbf{U}nequal \textbf{S}cores (AQUARIUS), for re-defining radiology QA, which reduces human effort by up to several orders of magnitude over existing approaches. AQUARIUS typically includes automatic comparison of AI-based image analysis with natural language processing (NLP) on radiology reports. Only the usually small subset of cases with discordant reads is subsequently reviewed by human experts. To demonstrate the clinical applicability of AQUARIUS, we performed a clinical QA study on Intracranial Hemorrhage (ICH) detection in 1936 head CT scans from a large academic hospital. Immediately following image acquisition, scans were automatically analyzed for ICH using a commercially available software (Aidoc, Tel Aviv, Israel). Cases rated positive for ICH by AI (ICH-AI+) were automatically flagged in radiologists' reading worklists, where flagging was randomly switched off with probability 50\%. Using AQUARIUS with NLP on final radiology reports and targeted expert neuroradiology review of only 29 discordantly classified cases reduced the human QA effort by 98.5\%, where we found a total of six non-reported true ICH+ cases, with radiologists' missed ICH detection rates of 0.52\% and 2.5\% for flagged and non-flagged cases, respectively. We conclude that AQUARIUS, by combining AI-based image analysis with NLP-based pre-selection of cases for targeted human expert review, can efficiently identify missed findings in radiology studies and significantly expedite radiology QA programs in a hybrid human-machine interoperability approach. 
\end{abstract}

% Include a list of keywords after the abstract 
\keywords{Artificial intelligence, quality assurance, radiology, intracranial hemorrhage, natural language processing, AQUARIUS }

\section{MOTIVATION} \label{sec:intro}  
The provision of high-quality health care is the goal of all medical services. In the case of radiology quality assurance (QA) programs, patient selection, the conduct of the examination, and the interpretation of the results can all have an impact on the achievement of this goal, such as by improving the diagnostic information content, reducing radiation exposure, reducing medical costs, and improving departmental management [\citeonline{1_SPIE22_davies1983quality}]. QA programs thus contribute to the provision of high-quality health care. 
 
Specifically, QA programs in radiology are of immense value in improving patient safety, optimizing the clinical usefulness of medical imaging exams, and increasing financial profits of radiology practices by lowering operational costs. Yet, contemporary approaches to radiology QA, such as widely used methods of randomly selecting cases for human review, suffer from limited effectiveness and efficiency. Hence, there is an urgent need for streamlining radiology QA programs to make them better and faster. Here, we present a novel hybrid human-machine interoperability approach, \textbf{A}rtificial Intelligence (AI)-Based \textbf{Qu}ality \textbf{A}ssurance by \textbf{R}estricted \textbf{I}nvestigation of \textbf{U}nequal \textbf{S}cores (AQUARIUS), for re-defining radiology QA, which has the potential of reducing human QA effort by several orders of magnitude over existing approaches.  
 
In the following, we will explain the basic idea of AQUARIUS, describe its clinical application within a retrospective clinical trial for QA in AI-based detection of Intracranial Hemorrhage (ICH) in emergent care head CT scans, and will report initial radiology QA results for this clinical application.  
\newline
This work is embedded in our group’s endeavor to expedite artificial intelligence in biomedical imaging by means of advanced pattern recognition and machine learning methods for computational radiology and radiomics, e.g., [\citeonline{
12_wismuller2006exploratory,
13_wismuller1998neural,
14_wismuller2002deformable,
15_behrends2003segmentation,
16_wismuller1997neural,
17_bunte2010exploratory,
18_wismuller1998deformable,
19_wismuller2009exploration,
20_wismuller2009method,
22_huber2010classification,
23_wismuller2009exploration,
24_bunte2011neighbor,
25_meyer2004model,
wismuller2001exploration,
saalbach2005hyperbolic,
wismuller2009computational,
26_wismuller2009computational,
leinsinger2003volumetric,
27_meyer2003topographic,
wismuller2009exploration,
28_meyer2009small,
29_wismueller2010model,
meyer2007unsupervised,
wismuller2000neural,
meyer2007analysis,
32_wismueller2008human,
wismuller2000hierarchical,
wismuller2015method,
33_huber2012texture,
37_otto2003model,
38_varini2004breast,
48_meyer2004computer,
40_meyer2004stability,
wismuller1998hierarchical,
41_meyer2008computer,
wismuller2013introducing,
45_bhole20143d,
46_nagarajan2013computer,
wismuller2001automatic,
wismueller1999adaptive,
meyer2005computer,
49_nagarajan2014computer,
pester2013exploring,
50_nagarajan2014classification,
yang2014improving,
wismuller2014pair,
wang2014investigating,
51_wismuller2014framework,
meyer2004local,
schmidt2014impact,
nagarajan2015integrating,
wismuller2015nonlinear,
nagarajan2015characterizing,
abidin2015volumetric,
wismuller2016mutual,
abidin2016investigating,
52_schmidt2016multivariate,
abidin2017classification,
abidin2017using,
61_dsouza2017exploring,
55_abidin2018deep,
chockanathan2018resilient,
dsouza2018mutual,
abidin2019investigating,
abidin2020detecting,
wismuller2020prospective,
dsouza2020large,
vosoughi2020large,
vosoughi2021marijuana, 
vosoughi2021eusipco,
dsouza2021large, 
vosoughi2021_LSAGC,
vosoughi2021_lsNonlinear,
vosoughi2021schizophrenia
}].
\section{METHODOLOGIC CONCEPT }
\subsection{Shortcomings of Current Approaches to Radiology QA}
The potential of conventional approaches to radiology QA is limited and subject to manifold shortcomings. A typical scenario for contemporary QA-related analysis would be to randomly select cases for human expert review, where medical imaging studies are compared with the findings described in the associated radiology reports. Although the scope of the analysis can be restricted to specific imaging study subsets, such as originating from given pre-defined radiology practices, scanners, imaging modalities, radiology exam types, patient origin (inpatient, outpatient, or emergency department), specific referring physicians, etc., the principal challenge remains the same, namely a lack of focus by not pre-selecting cases with a high ‘pre-test’ probability of discordant interpretation based on image content. 
A classic example are ‘peer-review’ approaches, such as the American College of Radiology (ACR) RADPEERTM program [\citeonline{2_SPIE22}], which restricts QA to cases with existing prior imaging studies. This program starts from the assumption that if, during interpretation of a new examination, there are prior images of the same area of interest, the interpreting radiologist will typically form an opinion of the previous interpretation while interpreting the new study. If the opinion of the previous interpretation is scored, a peer review event has occurred, where the reviewer scores the previous interpretation using a standardized rating scale. This approach requires significant human effort, excludes studies with no prior exams, does not distinguish between high and low ‘pre-test’ probability for discordant reads, so that only a tiny subset of inaccurate interpretations may be detected, given the usually low probability of clinically relevant discordant interpretations. In practice, most of the reviewed studies will be rated as ACR category 1 signifying a concordant read. This involves the issue of vigilance fatigue for rare events, e.g., [\citeonline{3_SPIE22_freudenburg1988perceived}], as known from non-medical domains, such as policing [\citeonline{4_SPIE22_krause2012vigilance}] or airport security, where the underlying probability of alert-inducing events is usually low. Case selection is usually based on a random approach, such as automatically requesting peer review for a certain number of cases (typically two) at the beginning of a radiologist’s shift, which will preclude the majority of inaccurate interpretations from being detected in the peer review process. In addition, such a peer-review approach does not only require human intervention by the original peer-reviewing radiologist, but also significant time and effort by a second human reviewer that performs a secondary review of cases with peer-review scores indicating discordant reads between current and prior studies. 
In conclusion, despite requiring a significant amount of human expert intervention time and effort, the vast majority of inaccurate radiology interpretations will remain undetected under the currently used radiology QA approaches. Hence, there is an urgent need to make the radiology QA process better and faster, with the goal to both improve patient safety and reduce costs 
\subsection{Artificial Intelligence (AI)-Based Quality Assurance by Restricted Investigation of Unequal Scores (AQUARIUS)}
AQUARIUS consists of two steps: In a first step, AI-based case analysis is compared to a ‘gold standard’ given by an automatically retrieved estimate of human case assessment. In a second step, cases with “unequal scores”, i.e., discordant assessments between AI-based and human interpretation, are reviewed by human experts. A typical diagnostic radiology implementation of AQUARIUS would include automatic comparison of AI-based image analysis results with natural language processing (NLP) results on final radiology reports. Only the usually small subset of cases with discordant reads would then be manually reviewed by dedicated subspecialty-trained expert radiologists. By restricting human effort to only those cases with a high ‘pre-test’ probability for missed or overcalled findings, AQUARIUS optimizes the use of human and fiscal resources spent on radiology QA programs, thus demonstrating a significant potential for both ‘saving lives’ and ‘saving money’ simultaneously. 
We emphasize that multiple organizational and technical challenges have to be met simultaneously for implementing a successful AQUARIUS study design. These challenges include, but are not limited to, (i) acquiring Institutional Review Board approval for automatic data retreival and radiologists’ participation in such QA initiatives, (ii) aligning the interests of diverse stakeholders, including radiologists, hospital administrators, AI solution vendors, and regulatory bodies, such as the ACR, and (iii) significant interface programming efforts for automatically retrieving heterogeneous information from multiple radiology and hospital IT systems, such as AI-based image analysis results and NLP-based classification of radiology reports for relevant clinical imaging findings.

\section{CLINICAL APPLICATION: INTRACRANIAL HEMORRHAGE DETECTION TRIAL }

\paragraph{Clinical Relevance:} To demonstrate the applicability of our novel hybrid human-AI quality assurance (QA) approach using the AQUARIUS framework, we performed a study for evaluating radiologists' performance on accurately reporting intracranial hemorrhage (ICH) in emergent care setting head CT scans. Timely detection of ICH on medical imaging studies is critical, because delayed therapeutic interventions in emergency settings may be detrimental for patient outcome. For example, in hemorrhagic stroke, the American Heart Association (AHA)/American Stroke Association (ASA) 2018 guidelines for early management emphasize the time-dependent benefit of tissue plasminogen activator (tPA) therapy [\citeonline{5_SPIOE22_powers20182018}]. Patients with ischemic stroke must be identified as soon as possible so that tPA can be administered within 3-4.5 hours of symptom onset [\citeonline{5_SPIOE22_powers20182018}]. A non-contrast head CT scan, which must be negative for hemorrhage, is a required diagnostic test before tPA can be administered in a timely manner [\citeonline{6_SPIE22_lees2010time,7_SPIE22_walter2012diagnosis,8_SPIE22_katzan2004utilization}]. The clinical outcomes of patients receiving tPA within a designated time limit are closely linked to time of treatment [\citeonline{8_SPIE22_katzan2004utilization,9_SPIE22_mazighi2013impact}]. Besides hemorrhagic stroke, there are numerous other clinical conditions, in which ICH will lead to rapid increase of intracranial pressure, where early detection and treatment of ICH can significantly reduce patient morbidity and mortality.  
\paragraph{Data and Methods:} A total of 1936 consecutive non-contrast emergency-setting head CT scans from 2 CT scanners at a large academic hospital were prospectively acquired over 47 consecutive days. Immediately following image acquisition, scans were automatically analyzed for ICH using commercially available software (Aidoc, Tel Aviv, Israel). Cases rated positive for ICH by AI (ICH-AI+) were automatically flagged in radiologists' reading worklists, where flagging was randomly switched off with probability 50\%, see e.g. [\citeonline{10_SPIE22_wismuller2020prospective}]. For ICH-AI+ cases, radiologists' missed ICH detection rates (ratio of number of missed ICH cases and number of all true ICH-AI+ cases) was calculated and compared between flagged and non-flagged ICH-AI+ cases, where images of all ICH-AI+ cases with ICH-negative results detected by natural language processing (NLP) of final radiology reports (ICH-AI+NLP- cases) were re-analyzed by an experienced neuroradiologist to identify true ICH+ cases missed by original radiology readings.

\section{RESULTS}\label{sec:results}
Among all 1936 CT scans, 381 ICH-AI+ cases were found, of which 190 cases were flagged. A total of 29 ICH-AI+NLP- cases were found, where 6 had been reported ICH+ by the radiology report. Of the remaining non-reported 23 ICH-AI+ cases, neuroradiology expert review identified 6 non-reported true ICH+ cases, where 5 cases were non-flagged, and only 1 case was flagged. This yields radiologists' missed ICH detection rates of 0.52\% and 2.5\% for flagged and nonflagged cases, respectively.  
\section{CONCLUSIONS}\label{sec:conclusions}
We present a novel hybrid human-machine interoperability approach, \textbf{A}rtificial Intelligence (AI)-Based \textbf{Qu}ality \textbf{A}ssurance by \textbf{R}estricted \textbf{I}nvestigation of \textbf{U}nequal \textbf{S}cores (AQUARIUS), for re-defining radiology QA, which has the potential of reducing human QA effort by several orders of magnitude over existing approaches, making radiology QA programs better and faster. AQUARIUS consists of two steps: In a first step, AI-based case analysis is compared to a ‘gold standard’ given by an automatically retrieved estimate of human case assessment. In a second step, cases with “unequal scores”, i.e., discordant assessments between AI-based and human interpretation, are reviewed by human experts. In our initial clinical study on detecting QA for detecting Intracranial Hemorrhage (ICH) on emergent noncontrast head CT scans, our results suggest that flagging ICH-AI+ cases on radiologists' worklists may decrease the rate of missed true ICH+ cases, although this was not found statistically significant based on small numbers of discordantly reported cases. Yet, our novel method of human expert review of ICH-AI+NLP- cases successfully revealed true ICH+ cases missed by original radiology reports with minimum human effort, and can therefore provide valuable contributions to radiology QA programs as an efficient hybrid human-AI approach. We conclude that, by combining AI image analysis with NLP-based pre-selection of cases for targeted human expert review, our AQUARIUS approach can efficiently identify missed findings in radiology reports and thus expedite radiology QA programs.

\renewcommand{\baselinestretch}{0.5}
\footnotesize

\acknowledgments % equivalent to \section*{ACKNOWLEDGMENTS}       
 
This research was partially funded by the American College of Radiology (ACR) Innovation Award “AI-PROBE: A Novel Prospective Randomized Clinical Trial Approach for Investigating the Clinical Usefulness of Artificial Intelligence in Radiology” (PI: Axel Wism{\"u}ller) and an Ernest J. Del Monte Institute for Neuroscience Award from the Harry T. Mangurian Jr. Foundation (PI: Axel Wism{\"u}ller). This work was conducted as a Practice Quality Improvement (PQI) project related to American Board of Radiology (ABR) Maintenance of Certificate (MOC) for A.W.
\newpage
% References
\bibliography{report,refs_citations} % bibliography data in report.bib
\bibliographystyle{spiebib} % makes bibtex use spiebib.bst

\end{document}